OpenStudy (anonymous):

Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9

OpenStudy (ajprincess):

subtract the first eqn from the second eqn. Vat do u get when u do so?

OpenStudy (aliahmed):

x=12 y=-1

OpenStudy (aliahmed):

3y=9-x y=9-x/3 then sub in eq 1 x-2*(9-x/3)=14 then 3x-18+2x=42 x=60/5=12

OpenStudy (anonymous):

y=-1 x=12 subtractin 1from2 we get -5y=5 therefore y=1 puttin y=-1 in 1 x =14-2 x = 12